OpenStudy (anonymous):

Sequences: Find a formula for S_n where n is greater or equal to 1: 1,3,7,15,31......

OpenStudy (wasiqss):

easy it is pippa i expect u to solve this :P

OpenStudy (anonymous):

well i never learnt this in my life

OpenStudy (anonymous):

first time i am touching upon it lol

OpenStudy (anonymous):

i know there is (4)^n

OpenStudy (anonymous):

thatt i see plus the previous term

OpenStudy (wasiqss):

c in each sequence the difference is getting multiplied by 2 like there is diff of 2 between 1 and 3 diff of 4 btw 3 and 7 and so

OpenStudy (anonymous):

lol i was totally offf heheheh

OpenStudy (anonymous):

okk i see that

OpenStudy (wasiqss):

did u got the sequence now

OpenStudy (anonymous):

so like how wld i set up the formula?

Directrix (directrix):

The nth term is (2^n) - 1. Now, the sum ....:)

OpenStudy (anonymous):

y -1?

OpenStudy (wasiqss):

pippa n is the term for which we have to find like first term means n=1

OpenStudy (anonymous):

ya i get that

OpenStudy (anonymous):

still a drop confused lol

OpenStudy (wasiqss):

good dumbo :P

OpenStudy (anonymous):

Thanks smarty pants :D

OpenStudy (anonymous):

ghost can u explain to me

OpenStudy (anonymous):

hahaha ok i see it thanks ghost

Directrix (directrix):

The sum of the first n terms where n is a positve integer can be found by the partial sum formula 2^ (n+1) - ( n + 2). Check it, please.

OpenStudy (anonymous):

but the first one u gave made sense

Directrix (directrix):

That was for a particular term. Say, you wanted to know the 8th term in the sequence. It would be (2^n) - 1 evaluated fo n = 8. So, the 8th term is 2^8 - 1 = 255.

OpenStudy (anonymous):

right ok

OpenStudy (anonymous):

so y did u give me this other formula?

Directrix (directrix):

2^(8 +1) - (8+2) = 2^9 - 10 = 502.

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

like how am i supposed to figure that out on my own? is there a process?

OpenStudy (anonymous):

ohh but i dont think it wants the addition of all th eterms it just wants like to find liek n terms

Directrix (directrix):

The sum of the entire series does not exist. The partial sum formula does.Partial means for a finite number of terms.

OpenStudy (anonymous):

i dont think they r referring to that fance formula i think they only want me to give the simple one

Directrix (directrix):

Now, the sum of the first 8 terms (1 + 3 + 7 + 15 + 31 + 63 + 127 + 255 = 502) should be the partial sum formula *** 2^ (n+1) - ( n + 2) *** evaluated for 8.

OpenStudy (anonymous):

ya but i dont think i need the sum.

Directrix (directrix):

Find a formula for S_n. S sub n *means* the sum of the first n terms, I think. I had an arithmetic error in my formula check. All's well.

OpenStudy (anonymous):

thanks ghost,i have anotehr one of these but will post it on anotehr thread

Directrix (directrix):

Okay.